6 ms·
Isn't it sad that since the NSA revelations anything is really "on the table" for our imaginations? Nothing seems too far-fetched these days...
by vxxzy 13y ago
Isn't it sad that since the NSA revelations anything is really "on the table" for our imaginations? Nothing seems too far-fetched these days...
- frou_dh 13y agoI know it's a hot topic, but the Snowden leak is hardly the genesis of distrust.
- jetti 13y agoFor me, it is more the case of Michael Hastings[1] that would bring foul play from a government agent rather than the Snowden case. [1]http://www.foxnews.com/us/2013/06/24/journalist-michael-hastings-sends-chilling-e-mail-to-colleagues-before-death/ http://www.foxnews.com/us/2013/06/24/journalist-michael-hast...
- 00rion 13y agoIndeed. I can't believe more people don't know about this.
- lawnchair_larry 13y agoI can't believe you crazies found HN.
- codex 13y agoThe U.S. mortality rate is 0.008 per year. That means that in any gathering of 100 people (roughly the number of speakers at Black Hat), there is a 55% chance that at least one will die in the year before the event and a 6.5% chance they would die one month before the event. Over ten years of conferences, the odds that someone would die within one month of speaking is 49%. The under-40 crowd doesn't really appreciate this since death predominantly affects older people.
- conroe64 13y agoThat doesn't consider that the age and economic status of conference speakers isn't representative of the U.S. as a whole. Not that it's impossible for it to be an accident, but the odds are certainly less than 49%, given that the crowd is mostly urban professionals.
- codex 13y agoSorry, I should be clearer: I was calculating based on the general mortality rate. The accident rate is much lower (0.000391). I agree that either measure is skewed, though; this is just a rough calculation. The annual death rate for 35-44 from all causes (see http://www.data360.org/dsg.aspx?Data_Set_Group_Id=587 http://www.data360.org/dsg.aspx?Data_Set_Group_Id=587) is 0.002, which puts the odds of a speaker death within one month of the conference at 15% over ten years. Poorer, but then again, this is for only one specific security conference. If you assume five security conferences a year of all sorts, it's back to 50% over ten years. I think it's a bit higher as these speakers live sedentary lifestyles, even for the U.S.
- tomjen3 13y agoYeah but that is for all causes and that includes things like long term cancer and other deceases that don't really strike suddenly.
- conroe64 13y agoMakes sense. Another thought: Barnaby Jack was one of the top speakers and was to speak on a very controversial subject. I would guess that out of the 5000 speeches that were presented in your scenario only a few of them, maybe 2%, would contain information controversial enough that foul play would appear as a reasonable scenario to an outside observer (and this is being generous). Let F = foul play occurred in order to disrupt a conference, D = death of speaker one month before conference Let P(D) = (.02 deaths per year for 25-35 y.o) / (12 months in a year) = 0.0017 Let P(F|D) = .001 (assuming 1 in 1000 chance foul play was involved given a death of a speaker at a conference) P(F) = P(F|^D) * P(^D) + P(F|D) * P(D) = 0 * 0.98 + 0.001 * .02 = 0.00002 Let P(D|F) = 1 (chance of death if foul play is involved, assumed 100%) So Bayes theorem gives us: P(F|D) = P(D|F)*P(F)/P(D) = 1 * 0.00002 / 0.0017 = 0.018 (chance of foul play for a single speech given the speaker died 1 month before) Let P(C) = 0.02 (probability of a controversial speech) P(D) = .0017 (from above) Let P(C&D|F) = .5 (assuming there is a 50% chance the speech was controversial given foul play did occur, and death always occurs from foul play) P(C&D) = .02 * 0.0017 = 0.000034 P(F) = 0.00002 (from above) P(F|C&D) = P(C&D|F) * P(F) / P(C&D) = 0.5 * 0.00002 / .000034 = around a 30% foul play was involved in Barnaby Jack's death There are a lot of assumptions here that could adjust the final figure up or down, but if I did my math right, foul play does seem a reasonable scenario, (but not a foregone conclusion). edit: removed line "P(F|D) = 0.00058 (from above)" as pointed out by user 0003. End result didn't change, though.
- nikster 13y agoIt is a little sad. Those days where one could rule out foul play by the US government in a case like this - those days are most certainly over. That said in this case we'll have to trust his family.